OpenStudy (anonymous):

10=4(3a+2) One more help please :P And this time the answer it gotta be a fraction. Its either gonna be 2/3 or 1/6, SOS :P

OpenStudy (anonymous):

a=1/6

OpenStudy (anonymous):

OK but can you like show works please?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Start by distributing the 4 over to 3a and 2.... 10=12a+8 subtract the 8 2=12a divide by 12 from both sides a=2/12 or 1/6

OpenStudy (anonymous):

OH ok i get it now thanks zordoloom... or whatever your real name is :)

OpenStudy (anonymous):

lol. Sure thing.

OpenStudy (anonymous):

Kk